OpenStudy (anonymous):

find midpoint of each line segment whose endpoints are given. a. (3,7), (7,7) b.(0,-6),(-4,2) c.(-5,-3), (-7,-1)

OpenStudy (anonymous):

distance formula almost like slope you take the x's add them togather then divide that by 2 and that's the point for x. You would do the same thing for y's

OpenStudy (anonymous):

example 3+7= 10/2= 5 7+7= 14/2= 7 so A would b (5,7)

OpenStudy (anonymous):

k, i get it but how would i do (7,5),(5-9)

OpenStudy (amistre64):

c = (-6,-2)

OpenStudy (amistre64):

6 is halfway between 7 and 5; -9 to 5 = 14 spots, 5-7 = -2 6,-2 is mid of 7,5 5,-9

OpenStudy (anonymous):

o ok great thanks i will try the rest...

OpenStudy (amistre64):

just keep in mind, mid point means halfway between the two.; find the middle of the ys and the middle of the xs...

OpenStudy (anonymous):

k:)

OpenStudy (anonymous):

thank u also marie92